# Sums of four fourth power of primes

### v2

@@ 3. Some Lemmas
$\begin{split}\prod_{\overset{13\leqslant p<N^{(1-\varepsilon)/32}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)&\leq\prod_{\overset{p\geqslant 13}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{7}}\right)\\ &<1\>+\>\sum_{n\geqslant 12}\>\frac{1}{n^{7}}<1\>+\>12^{-6},\end{split}$
@@
$|E-E^{*}|\leq(3\sqrt{p}+1)^{6}(p-1)(3p+1).$
then for $p\geq 101$, by $1+x\leq e^{x}$, we have
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{8}}}=e^{\sum_{p\geq 101}{\frac{E^{*}-E}{\left(p-1\right)^{8}}}},$
where
$\sum_{p\geq 101}{\frac{|E^{*}-E|}{\left(p-1\right)^{8}}}\leq 0.305,$
thus
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 1.357.$
In addition,
$\prod_{\overset{13\leq p\leq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 11.773.$
@@
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)<15.976$
@@
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<308e^{-\gamma}(1+\varepsilon)L^{-1}.$
It is not difficult to find that
$\frac{|E^{*}|}{p(p-1)^{7}}<\frac{1}{p^{2}}$
for $p$ large enough. Hence
$\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right)\leqslant\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
@@ 4. Proof of the Theorem
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6},$
@@
(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim V}1\sum_{p_{4}\sim V}1\geqslant(1-\varepsilon)\frac{U^{2}V^{2}}{\log^{2}U\log^{2}V}\\ &\geqslant\left(\frac{96}{5}\right)^{2}(1-\varepsilon)U^{2}V^{2}L^{-4}.\end{split}$
@@
(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bV^{4}L^{-8}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
@@
$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{96}{5}\biggr)^{4}U^{4}>\frac{1}{27241.64}N.$

### v5

@@ 3. Some Lemmas
$\begin{split}\prod_{\overset{17\leqslant p<N^{9(1-\varepsilon)/200}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)&\leq\prod_{\overset{p\geqslant 17}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{15}}\right)\\ &<1\>+\>\sum_{n\geqslant 19}\>\frac{1}{n^{15}}<1\>+\>19^{-14},\end{split}$
@@
$|E-E^{*}|\leq(3\sqrt{p}+1)^{14}(p-1)(3p+1).$
Now define $\delta_{p}$ as
$\left|\frac{E^{*}-E}{(p-1)^{16}}\right|\leq\frac{(3\sqrt{p}+1)^{14}(3p+1)}{(p-1)^{15}}=\delta_{p}.$
By numerical calculation
$\sum_{p\geq 17}\delta_{p}<0.03,$
then by $1+x\leq e^{x}$
$\prod_{\overset{17\leq p\leq N^{{9(1-\epsilon)}/{200}}}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)\leq\prod_{\overset{p\geq 17}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{16}}}\leq e^{\sum_{p\geq 17}{\delta_{p}}}\leq 1.1.$
@@
$\prod_{13\leqslant p<N^{(9(1-\varepsilon)/200}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)<1.1$
@@
$\prod_{17\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<231.713e^{-\gamma}(1+\varepsilon)L^{-1}.$
It is not difficult to find that for $p$ large enough.
$\frac{|E^{*}|}{p(p-1)^{15}}<\frac{1}{p^{2}}$
Hence
$\prod_{p\geqslant N^{9(1-\varepsilon)/200}}\left(1+\frac{E^{*}}{p(p-1)^{15}}\right)\leqslant\prod_{p\geqslant N^{9(1-\varepsilon)/200}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
@@ 4. Proof of the Theorem
$U=\left(\frac{N}{64(1+\delta_{0})}\right)^{1/4},\quad V=U^{7/8},$
@@
(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim U}1\sum_{p_{4}\sim U}1\sum_{p_{5}\sim V}1\sum_{p_{6}\sim V}1\sum_{p_{7}\sim V}\sum_{p_{8}\sim V}\\ &\geqslant(1-\varepsilon)\frac{U^{4}V^{4}}{\log^{4}U\log^{4}V}\geqslant\left(\frac{128}{7}\right)^{4}(1-\varepsilon)U^{4}V^{4}L^{-8}.\end{split}$
@@
(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bU^{4}V^{8}L^{-16}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
@@
$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{128}{7}\biggr)^{8}U^{4}>\frac{1}{414.465}N.$